### v1

@@ 4.1 Performance guarantee
Definition 4.1 ($\delta$-Offset-Expiry.).
A perishing process is $\delta$-offset-expiring if:
$\mathbb{P}\mathopen{}\mathclose{{}\left(\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\quad\forall\,t\geq 2}\right)\geq 1-\delta$
Theorem 4.2.
@@
  • $\textsc{Conf}^{P}_{t}(\eta_{t})=\frac{1}{2}\mathopen{}\mathclose{{}\left(\log(3t\log T/\delta)+\sqrt{\log^{2}(3t\log T/\delta)+8\eta_{t}\log(3t\log T/\delta)}}\right)$
Then, for any $\delta$-offset-expiring perishing process, with probability at least $1-3\delta$, Algorithm 1 achieves:
$\Delta_{\text{\it EF}}\lesssim\max\{L_{T},\mathcal{L}^{\textsf{perish}}+1/\sqrt{T}\}$ $\Delta_{\text{\it efficiency}}\lesssim\min\mathopen{}\mathclose{{}\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}}\right\}+T\mathcal{L}^{\textsf{perish}}$
@@
where $\lesssim$ drops poly-logarithmic factors of $T$, $\log(1/\delta)$, $o(1)$ terms, and absolute constants.
In Section B.2, we show that these bounds are indeed tight, relative to the lower bounds in Theorem 2.3 and Theorem 3.7. We dedicate the remainder of this section to further parsing the bounds on counterfactual envy and efficiency — graphically represented in Fig. 1 — given the scalings of $L_{T}$ and $\mathcal{L}^{\textsf{perish}}$. In Section 4.2 we derive necessary and sufficient conditions on the perishing process for $\delta$-offset-expiry to hold. We finally prove Theorem 4.2 in Section 4.3.
Note that our result is a strict generalization of Sinclair et al., (2022): in the simplest setting where $\mathcal{L}^{\textsf{perish}}=0$ (i.e., no perishing, or perishing with perfect predictions), we recover the trade-off they identified. The following corollary simplifies our bounds in the “low-envy” setting where $L_{T}\lesssim 1/\sqrt{T}$.
Corollary 4.3 (Low-Envy).
Suppose $L_{T}\lesssim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability at least $1-3\delta$:
@@
Corollary 4.3 implies that there is no efficiency benefit to increasing $L_{T}$ as long as $L_{T}\lesssim 1/\sqrt{T}$. When $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, our algorithm incurs $\widetilde{O}(\sqrt{T})$ envy and inefficiency. In this case, these quantities are driven by the exogenous uncertainty in demand. When $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$, on the other hand, envy and inefficiency are driven by unavoidable perishing due to prediction errors.
@@
Corollary 4.4 (High-Envy, High-Perishing).
Suppose $L_{T}\gtrsim 1/\sqrt{T}$ and $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability $1-3\delta$:
@@
Similarly in this regime, increasing $L_{T}$ doesn’t guarantee arbitrary gains in efficiency; thus, setting $L_{T}\sim\mathcal{L}^{\textsf{perish}}$ is optimal. We conclude by exploring the more nuanced “high-envy, low-perishing” regime. In this setting, our algorithm’s guarantees depend on whether the efficiency gain from increasing envy, $L_{T}^{-1}$, exceeds $T\mathcal{L}^{\textsf{perish}}$. We defer its proof to Section B.3.
Corollary 4.5 (High-Envy, Low-Perishing).
Suppose $L_{T}\gtrsim 1/\sqrt{T}$, $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, and $T\mathcal{L}^{\textsf{perish}}\lesssim L_{T}^{-1}$. Then, Perishing-Guardrail achieves with probability at least $1-3\delta$:
@@
On the other hand, if $T\mathcal{L}^{\textsf{perish}}\gtrsim L_{T}^{-1}$, Perishing-Guardrail achieves with probability at least $1-3\delta$:
@@
Thus, in “high-envy, low-perishing” regimes, if the cumulative $\sigma$-induced loss is order-wise dominated by the efficiency gains from envy (otherwise phrased, our allocation schedule $\sigma$ is high-enough quality that perishing is low), increasing $L_{T}$ allows Perishing-Guardrail to achieve inversely proportional gains in efficiency. One can do this until moving into the regime where $L_{T}^{-1}\lesssim T\mathcal{L}^{\textsf{perish}}$ (i.e., the cumulative $\sigma$-induced loss dominates efficiency gains from envy). At this point, further increasing $L_{T}$ hurts envy, and has no order-wise impact on efficiency.
Section 4.2 next provides conditions on the perishing distribution to satisfy $\delta$-offset expiry.
4.2 On $\delta$-offset expiry
In order to highlight the salient parameters of the perishing process that dictate offset-expiry, in this section we assume $B=N=T$ almost surely, with $N_{t}=1$ for all $t$. At the cost of cumbersome algebra, one can relax this assumption and derive entirely analogous results. In this case, $\delta$-offset expiry (Definition 4.1) reduces to $\mathbb{P}\mathopen{}\mathclose{{}\left(P_{<t}\leq t-1\ \forall\,t\geq 2}\right)\geq 1-\delta.$
For $t\in[T]$, let <$\mathcal{B}^{rand}_{<t}=\{b:\mathbb{P}\mathopen{}\mathclose{{}\left(T_{b}<t}\right)\in(0,1)\}$, and <$\mathcal{B}^{det}_{<t}=\{b:\mathbb{P}\mathopen{}\mathclose{{}\left(T_{b}<t}\right)=1\}$. Proposition 4.6 states that, in expectation, no more that $t-1$ items can perish for $\delta$-offset expiry to hold for non-trivial values of $\delta$. We defer all proofs in this section to Appendix B.4.
Proposition 4.6.
Suppose there exists $t\geq 2$ such that $\mathbb{E}[P_{<t}]>t-1$. If <$\mathcal{B}^{rand}_{<t}=\emptyset$, $\delta$-offset-expiry cannot be satisfied for any value of $\delta\in(0,1)$. Else, $\delta$-offset-expiry cannot be satisfied for $\delta<\frac{1}{2}-{\mathrm{Std}\mathopen{}\mathclose{{}\left[P_{<t}}\right]^{-3}}\cdot T$.
Note that this necessary condition fails to hold for one of the most standard models of perishing: geometrically distributed perishing with parameter $1/T$ (that is, a constant fraction of items perish in each period). This highlights that one of the most popular models in the literature is, in a sense, far too pessimistic; for this setting, there is no hope of achieving low envy and efficiency with high probability. Proposition 4.7 next establishes a sufficient condition for $\delta$-offset expiry to hold.
Proposition 4.7.
Suppose that $\mathbb{E}[P_{<t}]\leq t-1$ for all $t\geq 2$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq\sum_{t=2}^{T}\min\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\frac{\mathrm{Std}\mathopen{}\mathclose{{}\left[P_{<t}}\right]}{t-\mathbb{E}[P_{<t}]}}\right)^{2},\exp\mathopen{}\mathclose{{}\left(-\frac{{2}(t-\mathbb{E}[P_{<t}])^{2}}{|\mathcal{B}^{rand}_{<t}|}}\right)}\right\}\mathds{1}\{|\mathcal{B}^{rand}_{<t}|>0\}.$
Proposition 4.7 states that either the expected lag $t-\mathbb{E}[P_{<t}]$ must be large, or the coefficient of variation with respect to the random lag process $t-P_{<t}$, must be small. This then reduces to a bound on the variability of $P_{<t}$ in settings where perishing closely tracks demand. In our numerical experiments (Section 5) we instantiate the above bounds for common distributions.
@@ 4.3 Analysis
  • Defining and bounding the “good event”: (Section 4.3.1): We first show that, with probability at least $1-\delta$, the realizations of future arrivals and perishing are no worse than our algorithm’s pessimistic predictions. As a result, it suffices to condition over such “good” sample paths.
  • Establishing feasibility of $\underline{X}$ (Section 4.3.2): For this good event, we show that the static allocation $\underline{X}$ computed at the start of the horizon will never exhaust the budget, despite incurring $\sigma$-induced loss.
@@
4.3.1 Defining and bounding the “good event”
We analyze the performance of our algorithm under a so-called “good event” $\mathcal{E}$, the intersection of the following three events:
@@ 4.3.1 Defining and bounding the “good event”
    $\mathcal{E}_{N}=\mathopen{}\mathclose{{}\left\{|N_{(t,t^{\prime}]}-\mathbb{E}\mathopen{}\mathclose{{}\left[N_{(t,t^{\prime}]}}\right]|\leq\textsc{Conf}^{N}_{t,t^{\prime}}\ \forall\ t,t^{\prime}>t\,\,}\right\},$
@@
    $\mathcal{E}_{\overline{P}}=\mathopen{}\mathclose{{}\left\{\overline{P}_{t}\geq\textsc{PUA}^{alg}_{\geq t}\,\forall\ t\in\{0,1,\ldots,T\}}\right\}$, where $\textsc{PUA}^{alg}_{\geq t}$ denotes the quantity of unallocated items that perished between the end of round $t$ and the end of round $T-1$,
  3. 3.
    $\mathcal{E}_{oe}=\mathopen{}\mathclose{{}\left\{\frac{P_{<t}}{B}\leq\frac{N_{<t}}{{N}}\quad\forall\,t\geq 2}\right\}$.
$\mathcal{E}$ represents the event that the arrival process falls close to its mean, that $\overline{P}_{t}$ is indeed a pessimistic estimate of the unallocated goods that perish in the future, and that the process is offset-expiring.
Since the process is $\delta$-offset expiring, we have that $\mathbb{P}(\mathcal{E}_{oe})\geq 1-\delta$ by assumption. The following lemma provides the high-probability bound on $\mathbb{P}(\mathcal{E}_{N})$. We defer its proof — which follows from a standard application of Hoeffding’s inequality — to Appendix B.5.
Lemma 4.8.
@@
The main challenge in analyzing $\mathcal{E}$ lies in showing that $\overline{P}_{t}$ is indeed a pessimistic estimate of our algorithm’s future perishing. To upper bound the amount of unallocated resources that perished between $t$ and $T-1$, we must account for both the uncertainty in arrivals and the realized order in which resources perished, and relate these two sources of uncertainty to the time at which the algorithm intended to allocate these resources. Establishing this upper bound hinges upon the careful construction of the “slow” consumption process, which decouples future perishing from future allocations to compute $\overline{P}_{t}$. We formalize these ideas in Lemma 4.9.
Lemma 4.9.
@@
Proof of Lemma 4.9.
We prove the claim by induction.
@@
Inductive step: $t-1\rightarrow t.$ Since $\textsc{PUA}^{alg}_{\geq t}$ represents the amount of unallocated goods that perished between $t$ and $T-1$, we have:
$\textsc{PUA}^{alg}_{\geq t}$ ${=}\sum_{\tau=t}^{T-1}\sum_{b\in\mathcal{B}^{alg}_{\tau}}\mathds{1}\mathopen{}\mathclose{{}\left\{T_{b}=\tau,b\not\in\mathcal{A}_{\tau}}\right\}=\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\mathopen{}\mathclose{{}\left\{T_{b}\geq t,b\text{ not allocated before }T_{b}}\right\}$ (11)
Recall, ${\tau}_{b}(t\mid\underline{X},\sigma)$ (8) is an upper bound on the latest possible time the algorithm would have allocated $b$. This follows from the fact that the least the algorithm could have allocated before $t$ under $\mathcal{E}_{N}$ is $\underline{N}_{<t}\underline{X}$. Similarly, the least amount of goods that the algorithm can allocate between $t$ and $t^{\prime}\geq t$ is $\underline{N}_{[t,t^{\prime}]}\underline{X}$. Hence, if $\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)$ and $b$ did not perish before $t^{\prime}$, it must be that the algorithm allocated $b$. Applying this logic to (11), we have:
$\textsc{PUA}_{\geq t}^{alg}$ $\leq\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\mathopen{}\mathclose{{}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}}\right\},$ (12)
since it could have been that an item $b^{\prime}$ such that $\sigma(b^{\prime})<\sigma(b)$ perished early, resulting in an earlier allocation time for $b$. A similar argument gives us that $\mathcal{B}^{alg}_{t}\subseteq\overline{\mathcal{B}}_{t}$. Plugging this into (12):
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\sum_{b\in\overline{\mathcal{B}}_{t}}\mathds{1}\mathopen{}\mathclose{{}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}}\right\}.$ (13)
Recall, $\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\})$. Applying a Chernoff bound (see Corollary C.3) to the right-hand side of (13), we obtain that, with probability at least $1-\delta/(3t\log(T))$:
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\eta_{t}+\frac{1}{2}\mathopen{}\mathclose{{}\left(\log(3t\log(T)/\delta)+\sqrt{\log^{2}(3t\log(T)/\delta)+8\eta_{t}\log(3t\log(T)/\delta)}}\right)$
@@
Moreover, $\textsc{PUA}^{alg}_{\geq t}\leq\textsc{PUA}^{alg}_{\geq(t-1)}\leq\overline{P}_{t-1}$, where the second inequality follows from the inductive hypothesis. Putting these two facts together, we obtain:
$\textsc{PUA}^{alg}_{\geq t}\leq\min\mathopen{}\mathclose{{}\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})}\right\}=\overline{P}_{t}$
with probability at least $1-\delta/(3t\log(T))$. A union bound over $t$ completes the proof of the result. ∎
Lemma 4.10 follows from these high-probability bounds. We defer its proof to Section B.5.
Lemma 4.10.
Let $\mathcal{E}=\mathcal{E}_{N}\cap\mathcal{E}_{\overline{P}}\cap\mathcal{E}_{oe}$. Then, $\mathbb{P}(\mathcal{E})\geq 1-3\delta$.
@@ 4.3.2 Feasibility of $\underline{X}$
Proof of Lemma 4.11.
By induction on $t$.
@@
where the second inequality follows from the fact that $\overline{N}_{>t^{*}}\geq N_{>t^{*}}$ under $\mathcal{E}$. Thus, it suffices to show that $\overline{P}_{t^{*}}\geq\textsc{PUA}^{alg}_{[t^{*},t)}$. This holds since $\textsc{PUA}^{alg}_{[t^{*},t)}\leq\textsc{PUA}^{alg}_{\geq t^{*}}\leq\overline{P}_{t^{*}}$ by Lemma 4.9. ∎
4.3.3 Improving efficiency via $\overline{X}$
@@
Lemma 4.12.
Given $\mathcal{E}$, let $t_{0}=\sup\{t:X_{t}^{alg}=\underline{X}\}$ be the last time that $X_{t}^{alg}=\underline{X}$ (or else 00 if the algorithm always allocates according to $\overline{X}$). Then, for some $\tilde{c}=\widetilde{\Theta}(1)$,
$t_{0}>T-\tilde{c}\mathopen{}\mathclose{{}\left(\frac{1}{L_{T}}+\sqrt{\frac{T\mathcal{L}^{\textsf{perish}}}{L_{T}}}}\right)^{2}.$
We defer the proof of Lemma 4.12 to Appendix B.5. Observe that, as $\mathcal{L}^{\textsf{perish}}$ increases, our algorithm stops allocating $\underline{X}$ earlier on. We will next see that this loss propagates to the our final efficiency bound.
4.3.4 Putting it all together
@@
Proof of Theorem 4.2.
By Lemma 4.11, the algorithm never runs out of budget under event $\mathcal{E}$, which occurs with probability at least $1-3\delta$. As a result $X_{t,\theta}^{alg}\in\{\underline{X},\overline{X}\}$ for all $t\in[T]$, $\theta\in\Theta$. We use this to bound envy and efficiency.
@@
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ $\leq\frac{B}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]-\textsc{Conf}_{0,T}^{N}}-\frac{B}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]+\textsc{Conf}_{0,T}^{N}}$
$=\frac{B}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}\mathopen{}\mathclose{{}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}}\right)=\beta_{avg}\mathopen{}\mathclose{{}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}}\right).$
Using the fact that $\textsc{Conf}_{0,T}^{N}=\sqrt{2T|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}$ and $\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]=\Theta(T)$, there exists $c_{1},c_{2}=\widetilde{\Theta}(1)$ such that, for large enough $T$, $\mathopen{}\mathclose{{}\left(1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}\right)^{-1}\leq\mathopen{}\mathclose{{}\left(1-c_{1}/\sqrt{T}}\right)^{-1}\leq 1+2c_{1}/\sqrt{T}$ and $\mathopen{}\mathclose{{}\left(1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\mathopen{}\mathclose{{}\left[N}\right]}}\right)^{-1}\geq\mathopen{}\mathclose{{}\left(1+c_{2}/\sqrt{T}}\right)^{-1}\geq 1-c_{2}/\sqrt{T}$. Plugging this into the above:
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ $\leq\beta_{avg}\mathopen{}\mathclose{{}\left(1+2c_{1}/\sqrt{T}-(1-c_{2}/\sqrt{T})}\right)\leq\beta_{avg}(2c_{1}+c_{2})/\sqrt{T}\lesssim 1/\sqrt{T}.$

### v2

@@ 4.2 Performance Guarantee
Theorem 4.1 below provides upper bounds on counterfactual envy and inefficiency, given hindsight envy parameter $L_{T}$.
Theorem 4.1
@@
  • $\textsc{Conf}^{P}_{t}(\eta_{t})=\frac{1}{2}\left(\log(3t\log T/\delta)+\sqrt{\log^{2}(3t\log T/\delta)+8\eta_{t}\log(3t\log T/\delta)}\right)$
Then, with probability at least $1-2\delta$, Algorithm 1 achieves:
$\Delta_{\text{\it EF}}\lesssim\max\{L_{T},\mathcal{L}^{\textsf{perish}}+1/\sqrt{T}\}$ $\Delta_{\text{\it efficiency}}\lesssim\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\}+T\mathcal{L}^{\textsf{perish}}$
@@
where $\lesssim$ drops poly-logarithmic factors of $T$, $\log(1/\delta)$, and absolute constants.
In Section 8.1, we show that these bounds are tight relative to the lower bounds in Theorems 2.4 and 3.8. We dedicate the remainder of this section to further parsing the bounds on counterfactual envy and efficiency given the scalings of $L_{T}$ and $\mathcal{L}^{\textsf{perish}}$. We prove Theorem 4.1 in Section 4.3, and instantiate $\mathcal{L}^{\textsf{perish}}$ for important special cases in Section 4.4.
$L_{T}$$\mathcal{L}^{\textsf{perish}}$$T^{-1/2}$$T^{-1/2}$High-Envy,Low-Perishing(Corollary 4.4)Low-Envy(Corollary 4.2)High-Envy,High-Perishing(Corollary 4.3)
Figure 3: Graphical representation of the regimes described in Corollaries 4.2, 4.4 and 4.3. The curve $T\mathcal{L}^{\textsf{perish}}L_{T}=1$, which partitions the two cases of Corollary 4.4, is shown in gray.
The following corollary simplifies our bounds in the “low-envy” setting where $L_{T}\lesssim 1/\sqrt{T}$.
Corollary 4.2 (Low-Envy)
Suppose $L_{T}\lesssim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
@@
Corollary 4.2 implies that there is no efficiency benefit to increasing $L_{T}$ as long as $L_{T}\lesssim 1/\sqrt{T}$. When $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, our algorithm incurs $\widetilde{O}(\sqrt{T})$ envy and inefficiency. In this case, these quantities are driven by the exogenous uncertainty in demand. When $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$, on the other hand, envy and inefficiency are driven by the unavoidable perishing-induced loss.
@@
Corollary 4.3 (High-Envy, High-Perishing)
Suppose $L_{T}\gtrsim 1/\sqrt{T}$ and $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
@@
Similarly in this regime, increasing $L_{T}$ doesn’t guarantee arbitrary gains in efficiency; thus, setting $L_{T}\sim\mathcal{L}^{\textsf{perish}}$ is optimal. We conclude by exploring the more nuanced “high-envy, low-perishing” regime. In this setting, our algorithm’s guarantees depend on whether the efficiency gain from increasing envy, $L_{T}^{-1}$, exceeds $T\mathcal{L}^{\textsf{perish}}$. We defer its proof to Appendix 8.2.1.
Corollary 4.4 (High-Envy, Low-Perishing)
Suppose $L_{T}\gtrsim 1/\sqrt{T}$, $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, and $T\mathcal{L}^{\textsf{perish}}\lesssim L_{T}^{-1}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
@@
On the other hand, if $T\mathcal{L}^{\textsf{perish}}\gtrsim L_{T}^{-1}$, Perishing-Guardrail achieves with probability at least $1-2\delta$:
@@
Thus, in “high-envy, low-perishing” regimes, if the cumulative perishing-induced loss is order-wise dominated by the efficiency gains from envy (otherwise phrased, our allocation schedule $\sigma$ is high-enough quality and the perishing rate is mild), increasing $L_{T}$ allows Perishing-Guardrail to achieve inversely proportional gains in efficiency, as in the setting without perishable resources. One can do this until moving into the regime where $L_{T}^{-1}\lesssim T\mathcal{L}^{\textsf{perish}}$ (i.e., the cumulative perishing-induced loss dominates efficiency gains from envy). At this point, further increasing $L_{T}$ hurts envy, and has no order-wise impact on efficiency. We summarize these four regimes pictorially in Figure 3.
We conclude the section by noting that our result is a strict generalization of Sinclair et al. (2023): in the simplest setting where $\mathcal{L}^{\textsf{perish}}=0$ (i.e., no perishing, or the process is offset-expiring with perfect predictions), we recover the trade-off described in Theorem 2.4.
@@ 4.3 Analysis
  • Defining and bounding the “good event” (Section 4.3.1): We first show that, with probability at least $1-2\delta$, the realizations of future arrivals and perishing are no worse than our algorithm’s pessimistic predictions. As a result, it suffices to condition over such “good” sample paths.
  • Establishing feasibility of $\underline{X}$ (Section 4.3.2): For this good event, we show that the static allocation $\underline{X}$ computed at the start of the horizon will never exhaust the budget, despite incurring perishing-induced loss.
@@
4.3.1 Defining and bounding the “good event.”
We analyze the performance of our algorithm under a so-called “good event” $\mathcal{E}$, the intersection of the following two events:
@@
    $\mathcal{E}_{N}=\left\{|N_{(t,t^{\prime}]}-\mathbb{E}\left[N_{(t,t^{\prime}]}\right]|\leq\textsc{Conf}^{N}_{t,t^{\prime}}\ \forall\ t,t^{\prime}>t\right\},$
@@
    $\mathcal{E}_{\overline{P}}=\left\{\overline{P}_{t}\geq\textsc{PUA}^{alg}_{\geq t}\,\forall\ t\in\{0,1,\ldots,T\}\right\}$, where $\textsc{PUA}^{alg}_{\geq t}$ denotes the quantity of unallocated items that perished between the end of round $t$ and the end of round T1T-1.
$\mathcal{E}$ represents the event that the arrival process falls close to its mean, and that $\overline{P}_{t}$ is a pessimistic estimate of the unallocated goods that perish in the future.
The following lemma provides the high-probability bound on $\mathbb{P}(\mathcal{E}_{N})$. We defer its proof — which follows from a standard application of Hoeffding’s inequality — to Appendix 8.3.1.
Lemma 4.5
@@
The main challenge in analyzing $\mathcal{E}$ lies in showing that $\overline{P}_{t}$ is indeed a pessimistic estimate of our algorithm’s future perishing. To upper bound the amount of unallocated resources that perished between $t$ and T1T-1, we must account for both the uncertainty in arrivals and the realized order in which resources perished, and relate these two sources of uncertainty to the time at which the algorithm intended to allocate these resources. Establishing this upper bound hinges upon the careful construction of the “slow” consumption process. In particular, leveraging $\tau_{b}(t\mid\underline{X},\sigma)$ as an upper bound on the time that each unit is allocated by our algorithm, our bound on the number of units that perish from $t$ onwards is a sum of independent random variables that is decoupled from the algorithm’s past (and future) decisions, thereby allowing us to tractably obtain tight bounds on the true perishing process. We formalize these ideas in Lemma 4.6.
Lemma 4.6
@@
Proof 4.7
Proof of Lemma 4.6. We prove the claim by induction.
@@
Inductive step: $t-1\rightarrow t.$ Let $\mathcal{A}_{\tau}$ be the set of items allocated in round $\tau$, for $\tau\geq t$. Since $\textsc{PUA}^{alg}_{\geq t}$ represents the amount of unallocated goods that perished between $t$ and T1T-1, we have:
$\textsc{PUA}^{alg}_{\geq t}{=}\sum_{\tau=t}^{T-1}\sum_{b\in\mathcal{B}^{alg}_{\tau}}\mathds{1}\left\{T_{b}=\tau,b\not\in\mathcal{A}_{\tau}\right\}$ $=\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\left\{T_{b}\geq t,b\text{ not allocated before }T_{b}\right\}$
$\leq\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right\},$ (16)
where the inequality follows from the fact that $t\leq T_{b}<\min\{T,\tau_{b}(t\mid\underline{X},\sigma)\}$} is a necessary condition for item $b$ to have spoiled.
A similar argument gives us that $\mathcal{B}^{alg}_{t}\subseteq\overline{\mathcal{B}}_{t}$. Plugging this into (16):
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\sum_{b\in\overline{\mathcal{B}}_{t}}\mathds{1}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right\}.$ (17)
Recall, $\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\})$. Applying a Chernoff bound (see Corollary 10.3) to the right-hand side of (17), we obtain that, with probability at least $1-\delta/(3t\log(T))$:
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\eta_{t}+\frac{1}{2}\left(\log(3t\log(T)/\delta)+\sqrt{\log^{2}(3t\log(T)/\delta)+8\eta_{t}\log(3t\log(T)/\delta)}\right)$
@@
Moreover, $\textsc{PUA}^{alg}_{\geq t}\leq\textsc{PUA}^{alg}_{\geq t-1}\leq\overline{P}_{t-1}$, where the second inequality follows from the inductive hypothesis. Putting these two facts together, we obtain:
$\textsc{PUA}^{alg}_{\geq t}\leq\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}=\overline{P}_{t}$
[…]
Lemma 4.8 follows from these high-probability bounds. We defer its proof to Appendix 8.3.2.
Lemma 4.8
Let $\mathcal{E}=\mathcal{E}_{N}\cap\mathcal{E}_{\overline{P}}$. Then, $\mathbb{P}(\mathcal{E})\geq 1-2\delta$.
@@
Proof 4.10
Proof of Lemma 4.9. By induction on $t$.
@@
[…]
4.3.3 Improving efficiency via $\overline{X}$.
@@
Lemma 4.11
Given $\mathcal{E}$, let $t_{0}=\sup\{t:X_{t}^{alg}=\underline{X}\}$ be the last time that $X_{t}^{alg}=\underline{X}$ (or else 0 if the algorithm always allocates according to $\overline{X}$). Then, for some $\tilde{c}=\widetilde{\Theta}(1)$,
$t_{0}>T-\tilde{c}\left(\frac{1}{L_{T}}+\sqrt{\frac{T\mathcal{L}^{\textsf{perish}}}{L_{T}}}\right)^{2}.$
We defer the proof of Lemma 4.11 to Appendix 8.3.3. Observe that, as $\mathcal{L}^{\textsf{perish}}$ increases, our algorithm stops allocating $\underline{X}$ earlier on. We will next see how this loss propagates to our final efficiency bound.
4.3.4 Putting it all together.
@@
Proof 4.12
Proof of Theorem 4.1. By Lemma 4.9, the algorithm never runs out of budget under event $\mathcal{E}$, which occurs with probability at least $1-2\delta$. As a result $X_{t,\theta}^{alg}\in\{\underline{X},\overline{X}\}$ for all $t\in[T]$, $\theta\in\Theta$. We use this to bound envy and efficiency.
@@
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ ${=}\frac{B}{\mathbb{E}\left[N\right]-\textsc{Conf}_{0,T}^{N}}-\frac{B}{\mathbb{E}\left[N\right]+\textsc{Conf}_{0,T}^{N}}$
$=\frac{B}{\mathbb{E}\left[N\right]}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}\right)=\beta_{avg}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}\right).$
Using the fact that $\textsc{Conf}_{0,T}^{N}=\sqrt{2T|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}$ and $\mathbb{E}\left[N\right]=\Theta(T)$, there exists $c_{1},c_{2}=\widetilde{\Theta}(1)$ such that, for large enough $T$, $\left(1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}\right)^{-1}\leq\left(1-c_{1}/\sqrt{T}\right)^{-1}\leq 1+2c_{1}/\sqrt{T}$ and $\left(1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}\right)^{-1}\geq\left(1+c_{2}/\sqrt{T}\right)^{-1}\geq 1-c_{2}/\sqrt{T}$. Plugging this into the above:
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ $\leq\beta_{avg}\left(1+2c_{1}/\sqrt{T}-(1-c_{2}/\sqrt{T})\right)\leq\beta_{avg}(2c_{1}+c_{2})/\sqrt{T}\lesssim 1/\sqrt{T}.$